OpenStudy (anonymous):

what are the zeros of the polynomial function f(x)=(x+3)(x-1)(x+5) .....Are they 3, -1, and 5???/

OpenStudy (anonymous):

no they are -3,1, and -5

OpenStudy (anonymous):

-3,1,-5

OpenStudy (anonymous):

x =1,-5,-3

OpenStudy (anonymous):

THANK YOU YOU GUYS ROCK

OpenStudy (anonymous):

it is zero when you replace x by those numbers. for example if x is -3 you get \[(-3+3)(-3-1)(-3+5)=0\]

OpenStudy (anonymous):

Thanks for explaining !

OpenStudy (anonymous):

A zero is what you plug in to a variable to make the equation=0 so (-3+3), (-1+1) or (-5+5) would make that part of the equation=0, so when the rest of it is multiplied out it would still=0

OpenStudy (anonymous):

I have another one, f(x) = x^3 - x^2 - 6x?? What are the zeros for this??

OpenStudy (anonymous):

factor out an x first

OpenStudy (anonymous):

get \[x(x^2-x-6)\] then factor further as \[x(x-3)(x+2)\] so zeros are at \[\{0,-2,3\}\]

OpenStudy (anonymous):

-2, 0 and 3

OpenStudy (anonymous):

THANK YOU :) I AM A FAN

OpenStudy (anonymous):

Glad we could help :)

OpenStudy (anonymous):

1 073 741 824 * (4 * 4 * 4 * 4) = 274 877 906 944

OpenStudy (anonymous):

i used google and java